OpenStudy (anonymous):

help, medals will be given!! Solve y = 3x + 4 -2x – y = 21 What is the value of x?

OpenStudy (anonymous):

you could just sub 3x + 4 in for y in the 2nd equation to find x.

OpenStudy (studygurl14):

Are you doing substitution or elimination, or does it not matter? @maria.alison

OpenStudy (anonymous):

-2x - y = 21 -2x -(3x + 4) = 21 -2x -3x - 4 = 21 -5x = 21 + 4 -5x = 25 x = -5

OpenStudy (anonymous):

(-5,-11) like kelliegirl33 said, x=-5, and also, y=-11: put -5 in the y=3x+4 equation so y=3(-5)+4 y=-15+4 y=-11

OpenStudy (anonymous):

thanks yall!

OpenStudy (anonymous):

yw :)